5. (15 points) Find the line of intersection of the two planes. Show your work. 3x - 2y +1 2x+y - 3x = 3.
Answer #1


Given 37-24+2 = 1 20+ y -32=3 - 2x - 3x gives 6x-4y+22=2 6x + 3y -9 2=9 + -74 +112=-7 79-1127 Ty= 7+112 112+7 7 y 2 0 x 3 + g